OpenStudy (anonymous):

simplify tan^2 theta+1/tan^2 theta

OpenStudy (anonymous):

((tan^(2)(θ)+1)/(tan^(2)(θ)))=1+cot^(2)(θ)=csc^(2)(θ)

OpenStudy (anonymous):

thank you!!!!

OpenStudy (anonymous):

you are very welcome

OpenStudy (anonymous):

\[(\tan^2 (\theta)+1)/\tan^2(\theta)=1+\cot^2(\theta)=\csc^2(\theta)\]